in 6 years, B will be 2/3 as old as A; and 8 years ago, he was 1/5 as old.find their ages at present.
Answer by Jk22(389)

let a,b ages at present :

b+6=2/3(a+6)
b-8=1/5(a-8)

3b+18=2a+12
5b-40=a-8 -> a=5b-32, put in prec. : 3b+18=10b-64+12

7b=6+64=70 -> b=10, a=50-32=18.

Verif :

b+6=16, a+6=24, 24/3*2=8*2=16.
b-8=2, a-8=10, 10/5=2.